Diffusion Morning Report 04/12/12

Ted Berg tberg at stanford.edu
Thu Apr 12 06:18:15 PDT 2012


Good Morning All,
     Looking OK this morning. New issues: Silane for Epi and ICP is 
empty -new cylinder is on order. HCl for Epi tripped off again.
     Leftovers: Thermco Nitride-leak, POCl3 and BBr3 sensor issue, SiGe 
pump out for repair, No TLC on Thermco 3 and 4, Thermco 2 loader issue, 
TEOS needs a test. No other major issues at this time.

					
	Date: 	4/12/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*yellow* 	*tgo sensor issue-ran BBr3 run-observed* 	** 
** 	**
Tylan 6-POCl3 	*yellow* 	*tgo sensor issue tube reinstalled-ran POCl3 
run* 	** 	** 	**
Tylan 7-TEOS2 	*yellow* 	*needs test* 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*broken tube-bad pump-out for repair* 
*3.4/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*15.78/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*14.22/30* 	** 	**
Tylan 11-LTO 	*up* 	** 	*0./30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*down* 	*loader issue* 	** 	** 	*3*
Thermco 3 	*yellow* 	*testing TLC* 	** 	** 	**
Thermco4 	*yellow* 	*bad source controller-no TLC* 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*down* 	*leak* 	** 	** 	*2*
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*yellow* 	*hcl and silane issues* 	** 	** 	*1*
AMAT Epi Cham. B 	*down* 	*dopants on* 	** 	** 	**
TEL Spa 	*yellow* 	*needs test* 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	*changeover working well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20120412/342f6e34/attachment.html>


More information about the diffusion mailing list